OpenStudy (anonymous):

********************HELP**************************HELP**************** y varies directly as the square of x. When x = 1, y = 8. Find x when y = 128.

OpenStudy (anonymous):

finally someone to the rescue

OpenStudy (jdoe0001):

\(\bf y = (n)x^2\) if you were to solve for "n", what would "n" be?

OpenStudy (jdoe0001):

well, hold one

OpenStudy (anonymous):

n= y\[n=y\]

OpenStudy (anonymous):

n=y srt(x)?

OpenStudy (jdoe0001):

$$\bf y = (n)x^2\\ \text{when y=1, x=8, thus}\\ 1 = (n)8^2 $$ solve for "n" now :)

OpenStudy (anonymous):

1=(n)64 now what ?

OpenStudy (jdoe0001):

well yes, when something "varies directly to something meaning it moves at some proportion of something so as "x" changes value, "y" changes by some value

OpenStudy (jdoe0001):

$$ \bf y = (n)x^2\\ \text{when y=1, x=8, thus}\\ 1 = (n)8^2\\ \frac{1}{64} =n\\ \text{now what is "x" when y = 128}\\ 128 = (n)x \implies x = \cfrac{128}{n} $$

OpenStudy (jdoe0001):

hmm one sec, lemme revise that :/

OpenStudy (anonymous):

im still lost

OpenStudy (jdoe0001):

ok, what part ... you confuses you?

OpenStudy (anonymous):

why is that 128=(n)x

OpenStudy (anonymous):

i am so lost

OpenStudy (jdoe0001):

ohh, one sec -> Find x when y = 128. <--- the equation would be \( \bf y = (n)x^2\) so when y =128 the equation will look like \(\bf 128 = (n)x^2\)

OpenStudy (anonymous):

128 = 8*x²

OpenStudy (anonymous):

x² = 16 x = 4

OpenStudy (anonymous):

got it now

OpenStudy (jdoe0001):

$$\bf \text{now what is "x" when y = 128}\\ 128 = (n)x^2 \implies x = \sqrt{\cfrac{128}{n}} \implies x = \sqrt{\cfrac{128}{\frac{1}{64}}}\\ \implies \sqrt{\frac{128}{1} \times \frac{64}{1}} \implies x = \sqrt{8192} $$

OpenStudy (jdoe0001):

wait... smokes, I ... .messed up :/

OpenStudy (jdoe0001):

when x=1 y = 8

OpenStudy (jdoe0001):

so y = nx 8 = n1 n = 8

OpenStudy (jdoe0001):

so much for my simplifying

OpenStudy (anonymous):

lol

OpenStudy (jdoe0001):

heheh, but yes, you're correct :)

OpenStudy (jdoe0001):

I misread the terms values my bad